The volume of a cube is 64cm3. What is the total length of all the edges of the cube, in centimeters?
A)48 B) 32 C) 64 D) 40 E) 96

Answer by stanbon(75887)

The volume of a cube is 64cm3. What is the total length of all the edges of the cube, in centimeters?
-----
Volume = side^3
s^3 = 64
s = 4
--------
Each cube has 12 sides.
length of all the edges = 12*4 = 48 cm
=============================================
Cheers,
Stan H.

A)48 B) 32 C) 64 D) 40 E) 96